OpenStudy (anonymous):

Jonathon bought a pair of running shoes at 25% off and paid $69. What was the original price of the shoes? Please show all work.

OpenStudy (anonymous):

can somebody help me pls

OpenStudy (anonymous):

69/100=0.69 0.69*25=17.25 .69+17.2=86.25